# Robust $\Lambda$-quantiles and extreme probabilities

### v1

@@ 3 Robust $\Lambda$-quantiles
    $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(\bigwedge_{i=1}^{n}F_{i})$, $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(\bigvee_{i=1}^{n}F_{i})$, $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(\bigwedge_{i=1}^{n}F_{i})$, and $\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(\bigvee_{i=1}^{n}F_{i})$;
@@
    if $\Lambda$ is decreasing, then $\sup_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(\bigwedge_{i=1}^{n}F_{i})$, $\inf_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(\bigvee_{i=1}^{n}F_{i})$, $\sup_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(\bigwedge_{i=1}^{n}F_{i})$, and $\inf_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(\bigvee_{i=1}^{n}F_{i})$.
The property of attainability is crucial in the proof of (ii) of Theorem 1 and it is not redundant if the candidates in $\mathcal{M}$ is infinite. We have the following counterexample.
@@
$\Lambda(x)=\left\{\begin{array}[]{cc}3/4,&x<0\\ 1/2,&0\leqslant x<1\\ 1/4,&x\geqslant 1\end{array}\right.,$
@@
$F_{n}(x)=\left\{\begin{array}[]{cc}0,&x<0\\ 1/2+1/(2n),&0\leqslant x<1\\ 1,&x\geqslant 1\end{array}\right..$
@@
$F_{\mathcal{M}}^{-}(x)=\left\{\begin{array}[]{cc}0,&x<0\\ 1/2,&0\leqslant x<1\\ 1,&x\geqslant 1\end{array}\right.,$
@@
The above example shows that if the attainability fails to hold, then (ii) of Theorem 1 may not be valid. However, we still have the following conclusion:
@@
Hence, both $\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})$ and $q_{\Lambda}^{-}(F_{\mathcal{M}}^{+})$ are still useful by offering the upper and lower bounds of the corresponding robust $\Lambda$-quantiles.
Compared to the attainability of $F_{\mathcal{M}}^{+}$, the attainability of $F_{\mathcal{M}}^{-}$ may not be easy to check as it is difficult to employ the limit argument; see Section 4. Next, we provide an alternative assumption that is much easier to check if $\Lambda$ is decreasing. Let $\widehat{F}_{\mathcal{M}}^{-}(x)=\inf_{F\in\mathcal{M}}F(x-)$ for all $x\in\mathbb{R}$. We say $\widehat{F}_{\mathcal{M}}^{-}$ is attainable if for any $x\in\mathbb{R}$, there exists $F\in\mathcal{M}$ such that $F(x-)=\widehat{F}_{\mathcal{M}}^{-}(x)$.
@@
Based on Proposition 3, we next discuss the existence of the worst or best distribution for the robust $\Lambda$-quantiles.
@@
    If $\Lambda$ is decreasing and left-continuous, and $\widehat{F}_{\mathcal{M}}^{-}$ is continuous and attainable, and $x=\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})\in\mathbb{R}$, then there exists $F^{*}\in\mathcal{M}$ such that $\tilde{q}_{\Lambda}^{+}(F^{*})=\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)$;
@@
    If $\Lambda$ is right-continuous, $F_{\mathcal{M}}^{+}$ is attainable and right-continuous and $x=q_{\Lambda}^{-}(F_{\mathcal{M}}^{+})\in\mathbb{R}$, then there exists $F^{*}\in\mathcal{M}$ such that $q_{\Lambda}^{-}(F^{*})=\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)$.
@@
(i) In light of Proposition 3, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})=x\in\mathbb{R}$. By the definition, we have either $F_{\mathcal{M}}^{-}(x)\leqslant\Lambda(x)$ or there exists a sequence $x_{n}\uparrow x$ such that $F_{\mathcal{M}}^{-}(x_{n})\leqslant\Lambda(x_{n})$. Note that $F_{\mathcal{M}}^{-}=\widehat{F}_{\mathcal{M}}^{-}$. Hence, $F_{\mathcal{M}}^{-}$ is continuous. For the second case, using the continuity of $F_{\mathcal{M}}^{-}$ and the left-continuity of $\Lambda$, letting $n\to\infty$, we have $F_{\mathcal{M}}^{-}(x)\leqslant\Lambda(x)$. As $\widehat{F}_{\mathcal{M}}^{-}$ is attainable, there exists $F^{*}\in\mathcal{M}$ such that $\widehat{F}^{*}(x)=\widehat{F}_{\mathcal{M}}^{-}(x)=F_{\mathcal{M}}^{-}(x)$ with $\widehat{F}^{*}(x)=F^{*}(x-)$. This implies $\widehat{F}^{*}(x)\leqslant\Lambda(x)$. Hence, $\tilde{q}_{\Lambda}^{+}(\widehat{F}^{*})\geqslant x=\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)$. Note that the decrease of $\Lambda$ implies $\tilde{q}_{\Lambda}^{+}(F^{*})=\tilde{q}_{\Lambda}^{+}(\widehat{F}^{*})$. Hence, we have $\tilde{q}_{\Lambda}^{+}(F^{*})\geqslant\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)$, which implies $\tilde{q}_{\Lambda}^{+}(F^{*})=\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)$.
(ii) First, it follows from Theorem 1 that $\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(F_{\mathcal{M}}^{+})=x\in\mathbb{R}$. By definition, we have either $F_{\mathcal{M}}^{+}(x)\geqslant\Lambda(x)$ or there exists a sequence of $x_{n}\downarrow x$ such that $F_{\mathcal{M}}^{+}(x_{n})\geqslant\Lambda(x_{n})$. For the second case, it follows from the right-continuity of $F_{\mathcal{M}}^{+}$ and $\Lambda$ that $F_{\mathcal{M}}^{+}(x)\geqslant\Lambda(x)$. The attainability of $F_{\mathcal{M}}^{+}$ implies there exists $F^{*}\in\mathcal{M}$ such that $F^{*}(x)=F_{\mathcal{M}}^{+}(x)\geqslant\Lambda(x)$. This implies $q_{\Lambda}^{-}(F^{*})\leqslant x=\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)$. Hence, we have $q_{\Lambda}^{-}(F^{*})=\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)$. ∎
Note that for other $\Lambda$-quantiles, the best or worst case of distribution may not exist, even for the case when $\Lambda=\alpha$. For instance, for the uncertainty set characterized by moment constraints, the optimal distributions for $\sup_{F\in\mathcal{M}}q_{\alpha}^{-}(F)$ and $\inf_{F\in\mathcal{M}}q_{\alpha}^{+}(F)$ do not exist; see e.g., Corollary 4.1 of Bernard et al. (2024).
@@
    $\sup_{F\in\mathcal{M}}{q}_{\Lambda}^{-}(F)=\sup_{x\in\mathbb{R}}\{q_{\Lambda(x)}^{-}(F^{-}_{\mathcal{M}})\wedge x\},$ and $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)=\inf_{x\in\mathbb{R}}\{q_{\Lambda(x)}^{+}(F^{+}_{\mathcal{M}})\vee x\}$;
@@
    $\sup_{F\in\mathcal{M}}{q}_{\Lambda}^{+}(F)=\sup_{x\in\mathbb{R}}\{q_{\Lambda(x)}^{+}(F^{-}_{\mathcal{M}})\wedge x\}$ if $F_{\mathcal{M}}^{-}$ is attainable, and $\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)=\inf_{x\in\mathbb{R}}\{q_{\Lambda(x)}^{-}(F^{+}_{\mathcal{M}})\vee x\}$ if $F_{\mathcal{M}}^{+}$ is attainable.
@@ 4 Specific uncertainty sets
In this section, our attention is directed towards the uncertainty sets characterized by three distinct and widely used constraints: (a) moment constraints, (b) probability constraints via Wasserstein distance; and (c) marginal constraints in risk aggregation. Furthermore, our results are applied to the portfolio selection problem. Utilizing Theorem 1, we can clearly see that the robust $\Lambda$-quantiles can be derived based on the expressions of robust quantiles. Denote by $\mathcal{M}_{p}$ the set of all cdfs on $\mathbb{R}$ with finite $p$-th moment and $\mathcal{M}_{p}\left(\mathbb{R}^{n}\right)$ the set of all cdfs on $\mathbb{R}^{n}$ with finite $p$-th moment in each component.
@@ 4.1 Moment constraints
For any increasing function $g:(0,1)\to\mathbb{R}$, define $g_{+}^{-1}(x)=\inf\{t\in(0,1):g(t)>x\}$ with the convention that $\inf\emptyset=1$. In the following results, we derive the extreme probabilities over the moment uncertainty sets, i.e., $F_{\mathcal{M}_{p}(m,v)}^{-}$ and $F_{\mathcal{M}_{p}(m,v)}^{+}$.
@@
Moreover, in light of (i) of Corollary 1, we have $\sup_{F\in\mathcal{M}_{p}(m,v)}q_{\alpha}^{-}(F)=q_{\alpha}^{-}(F_{\mathcal{M}_{p}(m,v)}^{-})$. Hence, $q_{\alpha}^{-}(F_{\mathcal{M}_{p}(m,v)}^{-})=l(\alpha)$ for $\alpha\in(0,1)$. Note that $F_{\mathcal{M}_{p}(m,v)}^{-}(\cdot)$ is right-continuous. Hence, $F_{\mathcal{M}_{p}(m,v)}^{-}(x)=l_{+}^{-1}(x)$ for $x\in\mathbb{R}$. One can easily check that $l$ is strictly increasing and continuous over $(0,1)$. Thus we have that $F_{\mathcal{M}_{p}(m,v)}^{-}$ is continuous over $\mathbb{R}$, which further implies $\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}$ is continuous over $\mathbb{R}$.
@@
Moreover, it follows from (i) of Corollary 1, we have $\inf_{F\in\mathcal{M}_{p}(m,v)}q_{\alpha}^{+}(F)=q_{\alpha}^{+}(F_{\mathcal{M}_{p}(m,v)}^{+})$. Combining the above two equations, we have $q_{\alpha}^{+}(F_{\mathcal{M}_{p}(m,v)}^{+})=u(\alpha)$ for $\alpha\in(0,1)$. Note that $u(\alpha)$ is strictly increasing and continuous over $(0,1)$. Hence, $F_{\mathcal{M}_{p}(m,v)}^{+}(x)=u_{+}^{-1}(x)$ for $x\in\mathbb{R}$ and $F_{\mathcal{M}_{p}(m,v)}^{+}$ is continuous.
@@
Next, we show the attainability of $\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}$ and $F_{\mathcal{M}_{p}(m,v)}^{+}$. For any $x\in\mathbb{R}$, there exists $F_{k}\in\mathcal{M}_{p}(m,v)$ such that $\lim_{k\to\infty}F_{k}(x-)=\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}(x)$. By Chebyshev’s inequality, we have $\sup_{k\geqslant 1}(1-F_{k}(z))\leqslant\frac{v^{p}}{|z-\mu|^{p}}$ for $z>\mu$ and $\sup_{k\geqslant 1}F_{k}(z)\leqslant\frac{v^{p}}{|z-\mu|^{p}}$ for $z<\mu$. Hence, $\{F_{k},~{}k\geqslant 1\}$ is tight. It follows from Theorem 25.10 of Billingsley (1995)) that there exist a subsequence $\{F_{k_{l}},~{}l\geqslant 1\}$ and a distribution $F$ such that $F_{k_{l}}\to F$ weakly as $l\to\infty$. This implies $F_{k_{l}}^{-1}(t)\to F^{-1}(t)$ over $(0,1)$ a.e..
@@
Using the fact $\sup_{l\geqslant 1}\int_{0}^{1}|F_{k_{l}}^{-1}(t)|^{p}\mathrm{d}t\leqslant(v+|m|)^{p}$, we have for $y>0$
@@
For any $x\in\mathbb{R}$, there exists a sequence $F_{k}\in\mathcal{M}_{p}(m,v)$ such that $\lim_{k\to\infty}F_{k}(x)=F_{\mathcal{M}_{p}(m,v)}^{+}(x)$. Similarly as in the proof of the attainability of $\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}$, there exist a subsequence $\{F_{k_{l}},~{}l\geqslant 1\}$ and $F\in\mathcal{M}_{p}(m,v)$ such that $F_{k_{l}}\to F$ weakly as $l\to\infty$. Moreover, we could find $\varepsilon_{n}\downarrow 0$ such that $F$ is continuous at all $x+\varepsilon_{n}$. Using the fact that $F_{k_{l}}\to F$ weakly as $l\to\infty$, we have $F(x+\varepsilon_{n})=\lim_{l\to\infty}F_{k_{l}}(x+\varepsilon_{n})\geqslant\lim_{l\to\infty}F_{k_{l}}(x)=F_{\mathcal{M}_{p}(m,v)}^{+}(x)$. Letting $n\to\infty$, we obtain $F(x)\geqslant F_{\mathcal{M}_{p}(m,v)}^{+}(x)$, which implies $F(x)=F_{\mathcal{M}_{p}(m,v)}^{+}(x)$. Hence, $F_{\mathcal{M}_{p}(m,v)}^{+}$ is attainable. ∎
@@
    $\sup_{F\in\mathcal{M}_{p}(m,v)}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}_{p}(m,v)}^{-})$; $\inf_{F\in\mathcal{M}_{p}(m,v)}q_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(F_{\mathcal{M}_{p}(m,v)}^{+})$; $\inf_{F\in\mathcal{M}_{p}(m,v)}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(F_{\mathcal{M}_{p}(m,v)}^{+})$;
@@
For the case $p=2$, we obtain the explicit formulas for $F^{-}_{\mathcal{M}_{2}(m,v)}$ and $F^{+}_{\mathcal{M}_{2}(m,v)}$. For general case $p\neq 2$, the values of $F^{-}_{\mathcal{M}_{(p,m,v)}}$ and $F^{+}_{\mathcal{M}_{(p,m,v)}}$ can be computed numerically using (5). For the particular case that $\Lambda$ is decreasing, it is more convenient to compute the robust $\Lambda$-quantiles using the expressions in Proposition 5 since we do not need to compute the inverse functions $l$ and $u$ to obtain $F_{\mathcal{M}_{p}{(m,v)}}^{-}$ and $F_{\mathcal{M}_{p}{(m,v)}}^{+}$.
In the following example, we compare the results for normal, exponential and uniform distributions with the worst or best values of $q_{\Lambda}^{-},q_{\Lambda}^{+},\tilde{q}_{\Lambda}^{-}$ and $\tilde{q}_{\Lambda}^{+}$, respectively.
@@
Best case 0.51 0.51 0.51 0.51
Worst case 5.36 5.36 5.36 5.36
@@
Best case 0.90 0.90 0.90 0.90
Worst case 3.00 3.00 3.00 3.00

### v2

@@ 3 Robust $\Lambda$-quantiles
    $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(\bigwedge_{i=1}^{n}F_{i})$, $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(\bigvee_{i=1}^{n}F_{i})$, $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(\bigwedge_{i=1}^{n}F_{i})$, and $\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(\bigvee_{i=1}^{n}F_{i})$.
@@
    If $\Lambda$ is decreasing, then we have $\sup_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(\bigwedge_{i=1}^{n}F_{i})$, $\inf_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(\bigvee_{i=1}^{n}F_{i})$, $\sup_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(\bigwedge_{i=1}^{n}F_{i})$, and $\inf_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(\bigvee_{i=1}^{n}F_{i})$.
The property of attainability may not hold if the candidates in $\mathcal{M}$ are infinite. We have the following counterexample.
@@
$\Lambda(x)=\left\{\begin{array}[]{cc}3/4,&x<0,\\ 1/2,&0\leqslant x<1,\\ 1/4,&x\geqslant 1,\end{array}\right.$
@@
$F_{n}(x)=\left\{\begin{array}[]{cc}0,&x<0,\\ 1/2+1/(2n),&0\leqslant x<1,\\ 1,&x\geqslant 1.\end{array}\right.$
@@
$F_{\mathcal{M}}^{-}(x)=\left\{\begin{array}[]{cc}0,&x<0,\\ 1/2,&0\leqslant x<1,\\ 1,&x\geqslant 1,\end{array}\right.$
@@
The above example shows that if the attainability fails to hold, then (ii) of Theorem 1 may not be valid. However, we still have the following conclusions:
@@
Hence, both $\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})$ and $q_{\Lambda}^{-}(F_{\mathcal{M}}^{+})$ are still useful because they offer the upper and lower bounds of the corresponding robust $\Lambda$-quantiles.
Compared with the attainability of $F_{\mathcal{M}}^{+}$, the attainability of $F_{\mathcal{M}}^{-}$ may not be easy to check as it is difficult to employ the limit argument; see the examples of the three uncertainty sets in Section 4. Next, we provide an alternative assumption that is much easier to check if $\Lambda$ is decreasing. Let $\widehat{F}_{\mathcal{M}}^{-}(x)=\inf_{F\in\mathcal{M}}F(x-)$ for all $x\in\mathbb{R}$. We say $\widehat{F}_{\mathcal{M}}^{-}$ is attainable if for any $x\in\mathbb{R}$, there exists $F\in\mathcal{M}$ such that $F(x-)=\widehat{F}_{\mathcal{M}}^{-}(x)$. The attainability of $\widehat{F}_{\mathcal{M}}^{-}$ for three specific uncertainty sets will be discussed in Propositions 6, 8 and 11 in Section 4.
With the attainability of $\widehat{F}_{\mathcal{M}}^{-}$ and the monotonicity of $\Lambda$, we obtain the first equality in (ii) of Theorem 1.
@@
Next, we discuss the existence of the worst or best distribution for the robust $\Lambda$-quantiles.
@@
    If $\Lambda$ is decreasing and left-continuous, and $\widehat{F}_{\mathcal{M}}^{-}$ is continuous and attainable, and $\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})\in\mathbb{R}$, then there exists $F^{*}\in\mathcal{M}$ such that $\tilde{q}_{\Lambda}^{+}(F^{*})=\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)$.
@@
    If $\Lambda$ is right-continuous, $F_{\mathcal{M}}^{+}$ is attainable and right-continuous and $q_{\Lambda}^{-}(F_{\mathcal{M}}^{+})\in\mathbb{R}$, then there exists $F^{*}\in\mathcal{M}$ such that $q_{\Lambda}^{-}(F^{*})=\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)$.
@@
(i) In light of Proposition 3, we have $\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)=\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})$. For the sake of simplicity, we denote $\tilde{q}_{\Lambda}^{+}(F_{\mathcal{M}}^{-})$ by $z$. By the definition, we have either $F_{\mathcal{M}}^{-}(z)\leqslant\Lambda(z)$ or there exists a sequence $x_{n}\uparrow z$ such that $F_{\mathcal{M}}^{-}(x_{n})\leqslant\Lambda(x_{n})$. Note that $F_{\mathcal{M}}^{-}=\widehat{F}_{\mathcal{M}}^{-}$. Hence, $F_{\mathcal{M}}^{-}$ is continuous. For the second case, using the continuity of $F_{\mathcal{M}}^{-}$ and the left-continuity of $\Lambda$, letting $n\to\infty$, we have $F_{\mathcal{M}}^{-}(z)\leqslant\Lambda(z)$. As $\widehat{F}_{\mathcal{M}}^{-}$ is attainable, there exists $F^{*}\in\mathcal{M}$ such that $\widehat{F}^{*}(z)=\widehat{F}_{\mathcal{M}}^{-}(z)=F_{\mathcal{M}}^{-}(z)$ with $\widehat{F}^{*}(z)=F^{*}(z-)$. This implies that $\widehat{F}^{*}(z)\leqslant\Lambda(z)$; hence, $\tilde{q}_{\Lambda}^{+}(\widehat{F}^{*})\geqslant z=\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)$. Note that the decrease of $\Lambda$ implies that $\tilde{q}_{\Lambda}^{+}(F^{*})=\tilde{q}_{\Lambda}^{+}(\widehat{F}^{*})$. Hence, we have $\tilde{q}_{\Lambda}^{+}(F^{*})\geqslant\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)$, which implies that $\tilde{q}_{\Lambda}^{+}(F^{*})=\sup_{F\in\mathcal{M}}\tilde{q}_{\Lambda}^{+}(F)$.
(ii) First, it follows from Theorem 1 that $\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(F_{\mathcal{M}}^{+})$. Let $z=q_{\Lambda}^{-}(F_{\mathcal{M}}^{+})$. By definition, we have either $F_{\mathcal{M}}^{+}(z)\geqslant\Lambda(z)$ or there exists a sequence of $x_{n}\downarrow z$ such that $F_{\mathcal{M}}^{+}(x_{n})\geqslant\Lambda(x_{n})$. For the second case, it follows from the right-continuity of $F_{\mathcal{M}}^{+}$ and $\Lambda$ that $F_{\mathcal{M}}^{+}(z)\geqslant\Lambda(z)$. The attainability of $F_{\mathcal{M}}^{+}$ implies $F^{*}\in\mathcal{M}$ exists such that $F^{*}(z)=F_{\mathcal{M}}^{+}(z)\geqslant\Lambda(z)$. This implies $q_{\Lambda}^{-}(F^{*})\leqslant z=\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)$. Hence, we have $q_{\Lambda}^{-}(F^{*})=\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)$. ∎
Note that for other $\Lambda$-quantiles, the best-case or worst-case distribution may not exist, even for the case when $\Lambda=\alpha$. For instance, for the uncertainty set characterized by moment constraints, the optimal distributions for $\sup_{F\in\mathcal{M}}q_{\alpha}^{-}(F)$ and $\inf_{F\in\mathcal{M}}q_{\alpha}^{+}(F)$ do not exist; see e.g., Corollary 4.1 of Bernard et al. (2024).
@@
    $\sup_{F\in\mathcal{M}}{q}_{\Lambda}^{-}(F)=\sup_{x\in\mathbb{R}}\{q_{\Lambda(x)}^{-}(F^{-}_{\mathcal{M}})\wedge x\},$ and $\inf_{F\in\mathcal{M}}q_{\Lambda}^{+}(F)=\inf_{x\in\mathbb{R}}\{q_{\Lambda(x)}^{+}(F^{+}_{\mathcal{M}})\vee x\}$.
@@
    If $F_{\mathcal{M}}^{-}$ is attainable, then $\sup_{F\in\mathcal{M}}{q}_{\Lambda}^{+}(F)=\sup_{x\in\mathbb{R}}\{q_{\Lambda(x)}^{+}(F^{-}_{\mathcal{M}})\wedge x\}$ , and if $F_{\mathcal{M}}^{+}$ is attainable, then $\inf_{F\in\mathcal{M}}q_{\Lambda}^{-}(F)=\inf_{x\in\mathbb{R}}\{q_{\Lambda(x)}^{-}(F^{+}_{\mathcal{M}})\vee x\}$.
@@ 4 Specific uncertainty sets
In this section, our attention is directed towards the uncertainty sets characterized by three distinct and widely used constraints: (a) moment constraints, (b) probability constraints via the Wasserstein distance; and (c) marginal constraints in risk aggregation. Furthermore, our results are applied to the portfolio selection problem. By using Theorem 1, we can clearly see that the robust $\Lambda$-quantiles can be derived based on the expressions of robust quantiles. Denote by $\mathcal{M}_{p}$ the set of all cdfs on $\mathbb{R}$ with a finite $p$-th moment and $\mathcal{M}_{p}\left(\mathbb{R}^{n}\right)$ the set of all cdfs on $\mathbb{R}^{n}$ with a finite $p$-th moment in each component.
@@ 4.1 Moment constraints
For an increasing function $g:(0,1)\to\mathbb{R}$, define $g_{+}^{-1}(x)=\inf\{t\in(0,1):g(t)>x\}$ with the convention that $\inf\emptyset=1$. In the following results, we derive the extremal distributions over the moment uncertainty sets, i.e., $F_{\mathcal{M}_{p}(m,v)}^{-}$ and $F_{\mathcal{M}_{p}(m,v)}^{+}$.
@@
Moreover, in light of (i) of Corollary 1, we have $\sup_{F\in\mathcal{M}_{p}(m,v)}q_{\alpha}^{-}(F)=q_{\alpha}^{-}(F_{\mathcal{M}_{p}(m,v)}^{-})$. Hence, $q_{\alpha}^{-}(F_{\mathcal{M}_{p}(m,v)}^{-})=l(\alpha)$ for $\alpha\in(0,1)$. Note that $F_{\mathcal{M}_{p}(m,v)}^{-}(\cdot)$ is right-continuous. Therefore, $F_{\mathcal{M}_{p}(m,v)}^{-}(x)=l_{+}^{-1}(x)$ for $x\in\mathbb{R}$. One can easily check that $l$ is strictly increasing and continuous over $(0,1)$. Thus, we have that $F_{\mathcal{M}_{p}(m,v)}^{-}$ is continuous over $\mathbb{R}$, which further implies that $\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}$ is continuous over $\mathbb{R}$.
@@
Moreover, it follows from (i) of Corollary 1, we have $\inf_{F\in\mathcal{M}_{p}(m,v)}q_{\alpha}^{+}(F)=q_{\alpha}^{+}(F_{\mathcal{M}_{p}(m,v)}^{+})$. By combining the above two equations, we have $q_{\alpha}^{+}(F_{\mathcal{M}_{p}(m,v)}^{+})=u(\alpha)$ for $\alpha\in(0,1)$. Note that $u(\alpha)$ is strictly increasing and continuous over $(0,1)$. Hence, $F_{\mathcal{M}_{p}(m,v)}^{+}(x)=u_{+}^{-1}(x)$ for $x\in\mathbb{R}$ and $F_{\mathcal{M}_{p}(m,v)}^{+}$ is continuous.
@@
Next, we show the attainability of $\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}$ and $F_{\mathcal{M}_{p}(m,v)}^{+}$. For any $x\in\mathbb{R}$, there exists $F_{k}\in\mathcal{M}_{p}(m,v)$ such that $\lim_{k\to\infty}F_{k}(x-)=\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}(x)$. By Chebyshev’s inequality, we have $\sup_{k\geqslant 1}(1-F_{k}(z))\leqslant\frac{v^{p}}{|z-\mu|^{p}}$ for $z>\mu$ and $\sup_{k\geqslant 1}F_{k}(z)\leqslant\frac{v^{p}}{|z-\mu|^{p}}$ for $z<\mu$. Hence, $\{F_{k},~{}k\geqslant 1\}$ is tight. It follows from Theorem 25.10 of Billingsley (1995)) that there exist a subsequence $\{F_{k_{l}},~{}l\geqslant 1\}$ and a distribution $F$ such that $F_{k_{l}}\to F$ weakly as $l\to\infty$. This implies that $F_{k_{l}}^{-1}(t)\to F^{-1}(t)$ over $(0,1)$ a.e..
@@
Given that $\sup_{l\geqslant 1}\int_{0}^{1}|F_{k_{l}}^{-1}(t)|^{p}\mathrm{d}t\leqslant(v+|m|)^{p}$, we have for $y>0$
@@
For any $x\in\mathbb{R}$, there exists a sequence $F_{k}\in\mathcal{M}_{p}(m,v)$ such that $\lim_{k\to\infty}F_{k}(x)=F_{\mathcal{M}_{p}(m,v)}^{+}(x)$. Similarly as in the proof of the attainability of $\widehat{F}_{\mathcal{M}_{p}(m,v)}^{-}$, there exists a subsequence $\{F_{k_{l}},~{}l\geqslant 1\}$ and $F\in\mathcal{M}_{p}(m,v)$ such that $F_{k_{l}}\to F$ weakly as $l\to\infty$. Moreover, we could find $\varepsilon_{n}\downarrow 0$ such that $F$ is continuous at all $x+\varepsilon_{n}$. Using the fact that $F_{k_{l}}\to F$ weakly as $l\to\infty$, we have $F(x+\varepsilon_{n})=\lim_{l\to\infty}F_{k_{l}}(x+\varepsilon_{n})\geqslant\lim_{l\to\infty}F_{k_{l}}(x)=F_{\mathcal{M}_{p}(m,v)}^{+}(x)$. Letting $n\to\infty$, we obtain $F(x)\geqslant F_{\mathcal{M}_{p}(m,v)}^{+}(x)$, which implies $F(x)=F_{\mathcal{M}_{p}(m,v)}^{+}(x)$. Hence, $F_{\mathcal{M}_{p}(m,v)}^{+}$ is attainable. ∎
@@
    $\sup_{F\in\mathcal{M}_{p}(m,v)}\tilde{q}_{\Lambda}^{-}(F)=\tilde{q}_{\Lambda}^{-}(F_{\mathcal{M}_{p}(m,v)}^{-})$, $\inf_{F\in\mathcal{M}_{p}(m,v)}q_{\Lambda}^{+}(F)=q_{\Lambda}^{+}(F_{\mathcal{M}_{p}(m,v)}^{+})$, and $\inf_{F\in\mathcal{M}_{p}(m,v)}q_{\Lambda}^{-}(F)=q_{\Lambda}^{-}(F_{\mathcal{M}_{p}(m,v)}^{+})$;
@@
For the case where $p=2$, we obtain the explicit formulas for $F^{-}_{\mathcal{M}_{2}(m,v)}$ and $F^{+}_{\mathcal{M}_{2}(m,v)}$. For general case where $p\neq 2$, the values of $F^{-}_{\mathcal{M}_{(p,m,v)}}$ and $F^{+}_{\mathcal{M}_{(p,m,v)}}$ can be computed numerically using (5). For the particular case in which $\Lambda$ is decreasing, it is more convenient to compute the robust $\Lambda$-quantiles using the expressions in Proposition 5 since we do not need to compute the inverse functions $l$ and $u$ to obtain $F_{\mathcal{M}_{p}{(m,v)}}^{-}$ and $F_{\mathcal{M}_{p}{(m,v)}}^{+}$.
In the following example, we compare the $\Lambda$-quantiles of normal, exponential and uniform distributions with their worst-case and best-case values over $\mathcal{M}_{2}(m,v)$, respectively.
@@
Best-case 0.51 0.51 0.51 0.51
Worst-case 5.36 5.36 5.36 5.36
@@
Best-case 0.90 0.90 0.90 0.90
Worst-case 3.00 3.00 3.00 3.00